A ball of mass m=8 kg is suspended from a string of length L=5m the ball revovles in constant speed of horizontal circle of radius r=3m calculate the speed


1
Expert's answer
2021-02-24T12:47:32-0500

Let's apply the Newton's Second Law of Motion in projections on axis "x" and "y":


"Tsin\theta=\dfrac{mv^2}{r},""Tcos\theta=mg."

Let's divide the first equation by the second one:


"tan\theta=\dfrac{v^2}{gr}."

From this formula we can find the speed of the ball:


"v=\sqrt{rgtan\theta}."

We can find angle "\theta" from the geometry:


"sin\theta=\dfrac{r}{L},""\theta=sin^{-1}(\dfrac{r}{L})=sin^{-1}(\dfrac{3\ m}{5\ m})=36.8^{\circ}."

Finally, we can calculate the speed of the ball:


"v=\sqrt{3\ m\cdot9.8\ \dfrac{m}{s^2}\cdot tan36.8^{\circ}}=4.69\ \dfrac{m}{s}."